# Dominant vertices and attractors’ landscape for Boolean networks

### v2

@@ 4. Clover networks with signed majority rule
    The distribution of periods along the attractors is heavily skewed toward short periods, and although the theoretical upper bound grows exponentially with $\ell$, the observed periods were always much smaller. This is consistent with the fact that the distribution of cycles in a clover network generates an induced local function $\Phi:\mathtt{B}_{\ell}\to\mathtt{B}$ effectively depending on a few first coordinates. Indeed, in all of our simulations, the maximal observed period was upper bounded by $2\ell$. The distribution of periods appears to depend on both the folding probability $p$ and the inhibition probability $q$, but most notably on $p$. The mean period $\langle P\rangle:=\frac{1}{N_{\mathcal{A}}}\sum_{\mathcal{C}}P(\mathcal{C})$, appears to be a decreasing function of $p$, and for each $p$ fixed, it would be a unimodal function of $q$ with maximum presumably around $q=1/2$.
@@
    0.3 4.6087 2.8288 2.2530
    0.6 5.2200 3.1201 2.5710
    0.9 4.2089 2.2182 1.4327
    Table 2. The mean period, $\langle P\rangle$, averaged over 500 simulations, as a function of the folding and inhibition probabilities.
@@
[…]
    Table 3. The reduction in size, $|\mathcal{C}_{\mathcal{F}}|/|\mathcal{C}_{F}|$, averaged over 500 simulations, as a function of the folding and inhibition probabilities. In the last line, the expected mean reduction $2^{-(N-\ell)}$.
@@
    0.3 2.0659 1.5564 1.0985
    0.6 2.0528 1.6683 1.0789
    0.9 1.9647 1.4788 1.0347
    $\ell$ 5.8700 4.0660 2.6680
    Table 4. Reduction in maximal transient length, $\Delta\tau_{\max}$, averaged over 500 simulations, as a function of the folding and inhibition probabilities. In the last line, the recurrence length averaged over the same 500 numerical experiments, as a function of the folding probability. In all these exploratory simulations we set $N=10$.
@@ 5. Concluding Remarks
Given a set of dominant vertices in a Boolean network (equivalent to a feedback vertex set if the in-degree is everywhere positive), we define an automata network on those vertices. In the case that the recurrence length and the cardinality of the dominant set are small, the induced automata network constitutes a reduction of the original Boolean network, which preserves all the periodic attractors while reducing the corresponding basins. This reduction depends only on the topology of the original network, and in principle could be further improved by taking into account the particularities of the regulatory rules. This is observed in the case of clover networks with signed majority rule, where the induced local rule depends only on a few first coordinates. A refinement of this reduction, by taking into account the description of the local rules, will be considered in future works.
@@
Our results suggest that, given the interactions’ structure, a maximal possible complexity of the dynamics is encoded in the induced automata network defined by the dominant vertices. An interesting program would consist of studying the landscape of attractors for induced automata networks of increasing structural complexity, which would be the simplest representatives of each class of asymptotically equivalent systems.
@@
Acknowledgements

### v3

@@ 4. Clover networks with signed majority rule
    The distribution of periods along the attractors is heavily skewed toward short periods, with a clearly non-normal shape characterized by a decay with a thick tail, and although the theoretical upper bound grows exponentially with $\ell$, the observed periods were always much smaller. This is consistent with the fact that the distribution of cycles in a clover network generates an induced local function $\Phi:\mathtt{B}_{\ell}\to\mathtt{B}$ effectively depending on a few first coordinates. Indeed, in all of our simulations, the maximal observed period was upper bounded by $2\ell$. The distribution of periods appears to depend on both the folding probability $p$ and the inhibition probability $q$, but most notably on $p$. The mean period $\langle P\rangle:=\frac{1}{N_{\mathcal{A}}}\sum_{\mathcal{C}}P(\mathcal{C})$, appears to be a decreasing function of $p$, and for each $p$ fixed, it would be a unimodal function of $q$ with maximum presumably around $q=1/2$.
@@
    0.3 7.3  [1.0,13.6] 4.5  [1.0,8.0] 2.5  [1.0,4.0]
    0.6 7.5  [1.3,13.6] 4.5  [1.0,8.0] 2.6  [1.3,4.0]
    0.9 6.1  [1.3,11.0] 3.6  [1.3,6.0] 1.3  [1.3,1.3]
    Table 2. Median of the mean period $\langle P\rangle$ (computed per realization) together with the empirical $q_{0.05},q_{0.95}]$ interval over 500 simulations, as a function of the folding and inhibition probabilities.
@@
    0.3 0.507  [0.015,1.000] 0.035  [0.007,0.062] 0.009  [0.003,0.015]
    0.6 0.382  [0.015,0.750] 0.035  [0.007,0.062] 0.009  [0.003,0.015]
    0.9 0.257  [0.015,0.500] 0.035  [0.007,0.062] 0.009  [0.003,0.015]
    $2^{-(N-\ell)}$ 0.057 $\pm$ 0.067 0.016 $\pm$ 0.012 0.006 $\pm$ 0.002
    Table 3. Median of the reduction factor $|\mathcal{C}_{\mathcal{F}}|/|\mathcal{C}_{F}|$ over 500 realizations, together with the empirical $q_{0.05},q_{0.95}]$ interval. The last line shows the theoretical mean reduction $2^{-(N-\ell)}$, with $\ell$ reported as mean $\pm$ standard deviation. The comparatively large dispersion induced in $2^{-(N-\ell)}$ is a direct consequence of its exponential dependence on $\ell$, and does not indicate statistical instability of the sample.
@@
    0.3 2.0  [0.0,4.0] 2.0  [1.0,3.0] 1.5  [1.0,2.0]
    0.6 2.1  [0.3,4.0] 2.0  [1.0,3.0] 1.5  [1.0,2.0]
    0.9 2.3  [0.6,4.0] 1.8  [1.0,2.6] 1.0  [1.0,1.0]
    $\ell$ 5.87 $\pm$ 1.71 4.06 $\pm$ 1.07 2.66 $\pm$ 0.67
    Table 4. Median of the reduction in maximal transient length $\Delta\tau_{\max}$ over 500 realizations, together with the empirical $q_{0.05},q_{0.95}]$ interval, as a function of the folding and inhibition probabilities. In the last line, the recurrence length $\ell$ is reported as mean $\pm$ standard deviation. All simulations were performed with $N=10$.
@@ 5. Concluding Remarks
Given a set of dominant vertices in a Boolean network under a synchronous update scheme, we induce an automata network on those vertices, which is asymptotically equivalent to the original one. In this setting, when the recurrence length and the cardinality of the dominant set are small, the induced automata network constitutes a reduction of the original Boolean network, which preserves all the periodic attractors while reducing the corresponding basins. This reduction depends only on the topology of the original network and is valid only in the case of synchronous updating.
The network reduction, when possible, could be further improved by taking into account the particularities of the regulatory rules. This is observed in the case of clover networks with signed majority rule, where the induced local rule depends only on a few first coordinates. A refinement of this reduction, by taking into account the description of the local rules, will be considered in future works. Since dominant sets of vertices make sense for asynchronous updates, similar results are expected for the asynchronous update schemes. Nevertheless, the extension to that framework would require a different treatment of dependency propagation. Such an extension is of particular interest due to its potential impact on the dynamical analysis of bio-inspired networks and related discrete dynamical systems, and is therefore left for future work.
The proposed reduction does not provide an a priori estimate of its size, since the cardinality of dominant sets depends entirely on the network topology. In the examples presented here, the reduction can be extreme and the dynamics of a network with $N$ vertices is fully captured by a single dominant vertex, yielding an effective complexity scaling of order $1/N$. Typical reduction factors in biologically inspired networks are currently under investigation.
@@
Our results suggest that, given the interactions’ structure, a maximal possible complexity of the dynamics is encoded in the induced network defined by the dominant vertices. An interesting program would consist of studying the landscape of attractors for induced networks of increasing structural complexity, which would be the simplest representatives of each class of asymptotically equivalent systems.
@@
Code Availability
The implementation used to generate clover networks and compute their full and reduced dynamics is publicly available at: https://github.com/arletteespana/clover_dynamics
Acknowledgments